I was wrong in this question and I don't know how to answer this... please help me..
what is the largest circle which can be drawn from a square sheet with a side of 81cm? which area is greater, the circle or the square sheet?
Please help...TY if you can help me

Answer by chana(69)

what is the largest circle which can be drawn from a square sheet with a side of 81cm? which area is greater, the circle or the square sheet?
sol :
the largest circle drawn in a square is by taking the diagonal as diameter of the circle .
now about circle, its dia = sq.root of(81^2 + 81^2) =114.5 cm
area = pi * r^2 = pi *(114.5/2)^2.-------- (1)
area of square = 81^2 -------- (2)
compare the answer of 1 & 2 you will get the answer.